# Overcoming Saturation in Density Ratio Estimation by Iterated Regularization

### v1

@@ E.4 Detailed Empirical Results
In this section, we add all result tables for the datasets described in the main paper. Table 3-13 show all results on the MiniDomainNet image datasets, Table 14-24 show all results on the Amazon Reviews text datasets, and Table 25-35 show all results on the HHAR sensory datasets. Each table shows the results for all the respective datasets given one domain adaptation method from AdaTime. Averaging the results in these tables over the scenario datasets and combining them leads to Table 2.
@@
Table 3: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by AdvSKM Liu and Xue [2021].
@@
Table 4: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by CDAN Long et al. [2018].
@@
Table 5: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by CMD [Zellinger et al., 2017].
@@
Table 6: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by CoDATS [Wilson et al., 2020].
@@
Table 7: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DANN [Ganin et al., 2016].
@@
Table 8: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DDC [Tzeng et al., 2014].
@@
Table 9: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DeepCoral [Sun et al., 2017].
@@
Table 10: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DIRT [Shu et al., 2018].
@@
Table 11: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DSAN [Zhu et al., 2021].
@@
Table 12: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by HoMM [Chen et al., 2020].
@@
Table 13: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by MMDA [Rahman et al., 2020].
@@
Table 14: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by AdvSKM [Liu and Xue, 2021].
@@
Table 15: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by CDAN [Long et al., 2018].
@@
Table 16: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by CMD [Zellinger et al., 2017].
@@
Table 17: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by CoDATS [Wilson et al., 2020].
@@
Table 18: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DANN [Ganin et al., 2016].
@@
Table 19: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DDC [Tzeng et al., 2014].
@@
Table 20: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by Deep Coral [Sun et al., 2017].
@@
Table 21: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DIRT [Shu et al., 2018].
@@
Table 22: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DSAN [Zhu et al., 2021].
@@
Table 23: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by HoMM [Chen et al., 2020].
@@
Table 24: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by MMDA [Rahman et al., 2020].
@@
Table 25: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by AdvSKM [Liu and Xue, 2021].
@@
Table 26: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by CDAN [Long et al., 2018].
@@
Table 27: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by CMD [Zellinger et al., 2017].
@@
Table 28: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by CoDATS [Wilson et al., 2020].
@@
Table 29: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DANN [Ganin et al., 2016]
@@
Table 30: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DDC [Tzeng et al., 2014].
@@
Table 31: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DeepCoral [Sun et al., 2017].
@@
Table 32: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DIRT [Shu et al., 2018].
@@
Table 33: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DSAN [Zhu et al., 2021].
@@
Table 34: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by HoMM [Chen et al., 2020].
@@
Table 35: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by MMDA [Rahman et al., 2020].

### v2

@@ E.4 Detailed Empirical Results
In this section, we add all result tables for the datasets described in the main paper. Table 4-14 show all results on the MiniDomainNet image datasets, Table 15-25 show all results on the Amazon Reviews text datasets, and Table 26-36 show all results on the HHAR sensory datasets. Each table shows the results for all the respective datasets given one domain adaptation method from AdaTime. Averaging the results in these tables over the scenario datasets and combining them leads to Table 2.
@@
Table 4: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by AdvSKM (Liu & Xue, 2021).
@@
Table 5: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by CDAN (Long et al., 2018).
@@
Table 6: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by CMD (Zellinger et al., 2017).
@@
Table 7: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by CoDATS (Wilson et al., 2020).
@@
Table 8: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DANN (Ganin et al., 2016).
@@
Table 9: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DDC (Tzeng et al., 2014).
@@
Table 10: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DeepCoral (Sun et al., 2017).
@@
Table 11: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DIRT (Shu et al., 2018).
@@
Table 12: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by DSAN (Zhu et al., 2021).
@@
Table 13: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by HoMM (Chen et al., 2020).
@@
Table 14: Mean and standard deviation (after $\pm$) of target classification accuracy on MiniDomainNet for the aggregation method in objective (9) with models computed by MMDA (Rahman et al., 2020).
@@
Table 15: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by AdvSKM (Liu & Xue, 2021).
@@
Table 16: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by CDAN (Long et al., 2018).
@@
Table 17: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by CMD (Zellinger et al., 2017).
@@
Table 18: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by CoDATS (Wilson et al., 2020).
@@
Table 19: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DANN (Ganin et al., 2016).
@@
Table 20: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DDC (Tzeng et al., 2014).
@@
Table 21: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by Deep Coral (Sun et al., 2017).
@@
Table 22: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DIRT (Shu et al., 2018).
@@
Table 23: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by DSAN (Zhu et al., 2021).
@@
Table 24: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by HoMM (Chen et al., 2020).
@@
Table 25: Mean and standard deviation (after $\pm$) of target classification accuracy on Amazon Reviews for the aggregation method in objective (9) with models computed by MMDA (Rahman et al., 2020).
@@
Table 26: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by AdvSKM (Liu & Xue, 2021).
@@
Table 27: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by CDAN (Long et al., 2018).
@@
Table 28: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by CMD (Zellinger et al., 2017).
@@
Table 29: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by CoDATS (Wilson et al., 2020).
@@
Table 30: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DANN (Ganin et al., 2016)
@@
Table 31: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DDC (Tzeng et al., 2014).
@@
Table 32: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DeepCoral (Sun et al., 2017).
@@
Table 33: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DIRT (Shu et al., 2018).
@@
Table 34: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by DSAN (Zhu et al., 2021).
@@
Table 35: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by HoMM (Chen et al., 2020).
@@
Table 36: Mean and standard deviation (after $\pm$) of target classification accuracy on HHAR for the aggregation method in objective (9) with models computed by MMDA (Rahman et al., 2020).